# Private Online Community Detection for Censored Block Models

### v1

@@ F.3 Lower bounding the packing number $m$ with respect to $\operatorname{err}\operatorname{rate}$:
where $p^{\prime}=2a\times\frac{\log(n)}{n}+a^{2}(2\zeta^{2}-2\zeta-1)\times\frac{\log^{2}(n)}{n^{2}}=2a\frac{\log(n)}{n}+\mathcal{O}\left(\frac{\log^{2}(n)}{n^{2}}\right)$ In step (a), we set $\beta=n^{-1}$ and $\eta=n^{-\Delta}$, where $\Delta>0$.
Appendix G Performance of the Private Online Change Detection Algorithm
In this section, we provide a detailed upper bound for the detection delay for the detection algorithm in Section 4. For simplicity, we perform the analysis on the data sequence $\{{\mathbf{A}}_{t},t\in\mathbb{N}\}$, and here ${\mathbf{A}}_{t}$ can stands for both the original data and the perturbed private data, since they both satisfy the CBM model. Denote the KL information number $I_{0}={\mathbb{E}}_{0}^{\bm{\sigma}^{\text{post}}}\left[\ell({\mathbf{A}}_{t};\bm{\sigma}^{\text{pre}},\bm{\sigma}^{\text{post}})\right]$
Average Run Length
We begin by considering the average run length, which characterizes average false alarm period when there is no change in the community structure.
Proof of Lemma 5.2.
We compute the average false alarm period using similar ideas as in Lemma 8.2.1 in [32]. For $t>w$ let us define a Shiryaev-Roberts like statistic $\{R_{t}\}$ as
$R_{t}=(R_{t-1}+1)e^{\ell({\mathbf{A}}_{t};\bm{\sigma}^{\text{pre}},\bm{\hat{\sigma}}_{t-1})},~{}R_{w}=0.$
@@ Appendix G Performance of the Private Online Change Detection Algorithm
${\mathbb{E}}_{\infty}[R_{t}-t|{\mathcal{F}}_{t-1}]={\mathbb{E}}_{\infty}\left[(R_{t-1}+1)e^{\ell({\mathbf{A}}_{t};\bm{\sigma}^{\text{pre}},\bm{\hat{\sigma}}_{t-1})}-t\Big{|}{\mathcal{F}}_{t-1}\right]=R_{t-1}-(t-1).$
The last equality is true because given ${\mathcal{F}}_{t-1}$ we have $\bm{\hat{\sigma}}_{t-1}$ fixed and is independent of the observation ${\mathbf{A}}_{t}$. Since $\ell({\mathbf{A}}_{t};\bm{\sigma}^{\text{pre}},\bm{\hat{\sigma}}_{t-1})$ is a valid log density ratio for ${\mathbf{A}}_{t}$, we have the its expectation under ${\mathbb{E}}_{\infty}$ regime equals to one. The martingale property of $\{R_{t}-t\}$ and usage of Optional Sampling allows us to write for any stopping time $T$ with finite expectation that
${\mathbb{E}}_{\infty}[R_{T}-T]={\mathbb{E}}_{\infty}[R_{w}-w]=-w~{}\Rightarrow~{}{\mathbb{E}}_{\infty}[T]-w={\mathbb{E}}_{\infty}[R_{T}].$ (69)
@@
$e^{S_{t}}=\max\{e^{S_{t-1}},1\}e^{\ell({\mathbf{A}}_{t};\bm{\sigma}^{\text{pre}},\bm{\hat{\sigma}}_{t-1})},~{}e^{S_{w}}=1.$
@@
Using induction, the fact that $e^{S_{w+1}}=R_{w+1}$ and that for $x\geq 0$, $x+1\geq\max\{x,1\}$, it is straightforward to prove that for $t>w$ we have $R_{t}\geq e^{S_{t}}$. With this observation and using (69) we can now write
$e^{b}\leq{\mathbb{E}}_{\infty}[e^{S_{T}}]\leq{\mathbb{E}}_{\infty}[R_{T}]={\mathbb{E}}_{\infty}[T]-w\leq{\mathbb{E}}_{\infty}[T],$
which proves the desired inequality. ∎
Given the above Lemma, it is obvious to see that we can choose the detection threshold
$b=\log\gamma$
such that the average run length constraint ${\mathbb{E}}_{\infty}[T]\geq\gamma$ is satisfied. Therefore, we adopt this threshold in subsequent delay analysis.
Detection Delay
In the following, we provide theoretical guarantees on the detection delay of the procedure (14). We first define the quantities, $\hat{I}_{0}={\mathbb{E}}_{0}^{\bm{\sigma}^{\text{post}}}\left[\ell({\mathbf{A}}_{t};\bm{\sigma}^{\text{pre}},\bm{\hat{\sigma}}_{t-1})\right]$, which resemble the KL information number $I_{0}$ above.
Theorem G.1 (WADD Upper Bound).
@@
Assume $\hat{I}_{0}>0$. Under the assumptions of Theorem 5.4, we have the following upper bound for the worst-case performance of the detection procedure (17),
${\mathbb{E}}_{0}^{\bm{\theta}^{\text{post}}}[T]\leq\frac{\log\gamma}{\hat{I}_{0}}(1+o(1)).$ (70)
Proof.
The proof follows from Theorem 1 of [36], after observing that $\log\gamma$ dominates the numerator term under our assumption. ∎
In the following, we analyze the empirical quantities $\hat{I}_{0}$ and establish the asymptotic optimality result. The key component in the analysis is the estimation error of the MLE $\hat{\bm{\sigma}}_{t}$. We first establish the probability of false estimation.
Lemma G.1 (Estimating $\bm{\sigma}$).
Under the post-change regime, if $a(\sqrt{1-\zeta}-\sqrt{\zeta})^{2}>1$, the estimate $\bm{\hat{\sigma}}$ resulted from $w$ i.i.d. post-change samples (adjacency matrices) satisfies, as $n\to\infty$,
$\operatorname{Pr}(\bm{\hat{\sigma}}_{t}=\bm{\sigma}^{\text{post}})\geq 1-wn^{-\Omega(1)}.$
Proof.
From the analysis in Section D, we write $\bar{\mathbf{A}}=\sum_{i=1}^{w}{\mathbf{A}}_{i}$. Following the result of Theorem $9$ in [17], for any constant $c>0$, there exists $c^{\prime}>0$ such that for any $n\geq 1$,
$\Pr\left\{\left\|\bar{\mathbf{A}}-{\mathbb{E}}[\bar{\mathbf{A}}]\right\|_{2}\leq c^{\prime}w\sqrt{np}\right\}\geq 1-wn^{-c}.$
Following the proof to Theorem 6 in [17], we have the desired result. ∎
Under Lemma G.1, we have
$\hat{I}_{0}=I_{0}\Pr\{\bm{\hat{\sigma}}_{t}=\bm{\sigma}^{\text{post}}\}+{\mathbb{E}}[\hat{I}_{0}|\bm{\hat{\sigma}}_{t}\neq\bm{\sigma}^{\text{post}}]\Pr\{\bm{\hat{\sigma}}_{t}\neq\bm{\sigma}^{\text{post}}\}$
@@
note that $\mathbb{E}[\hat{I}_{0}]=\mathbb{E}[\log\frac{p(A_{t};\hat{\bm{\sigma}}_{t})}{p(A_{t};\bm{\sigma}^{\text{pre}})}]=\mathbb{E}[\log\frac{p(A_{t};\hat{\bm{\sigma}}_{t})}{p(A_{t};\bm{\sigma}^{\text{post}})}+\log\frac{p(A_{t};\bm{\sigma}^{\text{post}})}{p(A_{t};\bm{\sigma}^{\text{pre}})}]=I_{0}-I(\hat{\bm{\sigma}}_{t},\bm{\sigma}^{\text{post}})$.
We note that when estimating the post-change community structure within the support that satisfying Assumption 1, we have $I(\hat{\bm{\sigma}}_{t},\bm{\sigma}^{\text{post}})\leq 2\Delta n\left(\log\frac{1-\zeta}{\zeta}\right)p(1-2\zeta)$ (see Lemma G.2 below, together with Remark 5.2), and thus $|I_{0}-I(\hat{\bm{\sigma}}_{t},\bm{\sigma}^{\text{post}})|\leq 3\Delta n\left(\log\frac{1-\zeta}{\zeta}\right)p(1-2\zeta)$. Note that
$\frac{\hat{I}_{0}}{I_{0}}=\frac{I_{0}\Pr\{\bm{\hat{\sigma}}_{t}=\bm{\sigma}^{\text{post}}\}+{\mathbb{E}}[\hat{I}_{0}|\bm{\hat{\sigma}}_{t}\neq\bm{\sigma}^{\text{post}}]\Pr\{\bm{\hat{\sigma}}_{t}\neq\bm{\sigma}^{\text{post}}\}}{I_{0}}=1-O(wn^{-c}),$
@@
and when combined with (70), we obtain the desired first-order optimality in Theorem 5.4.
Lemma G.2.
Under Assumption 1, we have
(${n\choose 2}-C_{\bm{\sigma}^{\text{pre}},\bm{\sigma}^{\text{post}}}\leq 2\Delta n,$ (71)
where $C_{\bm{\sigma},\bm{\sigma}^{\prime}}:=\sum_{i<j}\bm{\sigma}_{i}\bm{\sigma}_{j}\bm{\sigma^{\prime}}_{i}\bm{\sigma^{\prime}}_{j}$.
Indeed, when there are $\Delta$ users who change their membership between $+1$ and $-1$ after the change, we have the quantity (${n\choose 2}-C_{\bm{\sigma}^{\text{pre}},\bm{\sigma}^{\text{post}}}$ is upper bounded by $2\Delta(n-\Delta)$.
Appendix H Fundamental Limits of Private Online Change Detection

### v2

@@ Appendix C Proofs for Section IV-A
where $p^{\prime}=2a\times\frac{\log(n)}{n}+a^{2}(2\zeta^{2}-2\zeta-1)\times\frac{\log^{2}(n)}{n^{2}}=2a\frac{\log(n)}{n}+\mathcal{O}\left(\frac{\log^{2}(n)}{n^{2}}\right)$ . In step (a), we set $\beta=n^{-1}$ and $\eta=n^{-\Delta}$, where $\Delta>0$.
Appendix D Proofs for Section IV-B
[…]
D-A Proof of Lemma IV.1
We compute the average false alarm period using similar ideas as in Lemma 8.2.1 in [36]. Let us define a Shiryaev-Roberts like statistic $\{R_{t}\}$ as
$R_{t}=(R_{t-1}+1)e^{\tilde{\ell}(\tilde{\mathbf{A}}_{t};\bm{\sigma}^{\text{pre}},\bm{\hat{\sigma}}_{t-1})},t>1,\qquad~R_{1}=0.$
@@ Appendix D Proofs for Section IV-B
${\mathbb{E}}_{\infty}[R_{t}-t|{\mathcal{F}}_{t-1}]$ $={\mathbb{E}}_{\infty}\left[(R_{t-1}+1)e^{\tilde{\ell}(\tilde{\mathbf{A}}_{t};\bm{\sigma}^{\text{pre}},\bm{\hat{\sigma}}_{t-1})}-t\Big|{\mathcal{F}}_{t-1}\right]$
$=R_{t-1}-(t-1).$
The last equality is true because given ${\mathcal{F}}_{t-1}$ we have $\bm{\hat{\sigma}}_{t-1}$ fixed and is independent of the observation ${\mathbf{A}}_{t}$. Since $\tilde{\ell}(\tilde{\mathbf{A}}_{t};\bm{\sigma}^{\text{pre}},\bm{\hat{\sigma}}_{t-1})$ is a valid log density ratio for $\tilde{\mathbf{A}}_{t}$, we have that its expectation under $\mathbb{E}$$_{\infty}$ regime equals to one. The martingale property of $\{R_{t}-t\}$ and usage of Optional Sampling allows us to write for any stopping time $T$ with finite expectation that
${\mathbb{E}}_{\infty}[R_{T}-T]={\mathbb{E}}_{\infty}[R_{1}-1]=-1~\Rightarrow~{\mathbb{E}}_{\infty}[T]-1={\mathbb{E}}_{\infty}[R_{T}].$ (48)
@@
$e^{S_{t}}=\max\{e^{S_{t-1}},1\}e^{\tilde{\ell}(\tilde{\mathbf{A}}_{t};\bm{\sigma}^{\text{pre}},\bm{\hat{\sigma}}_{t-1})},~e^{S_{1}}=1.$
@@
Using induction, the fact that $e^{S_{2}}=R_{2}$ and that for $x\geq 0$, $x+1\geq\max\{x,1\}$, it is straightforward to prove that for $t>1$ we have $R_{t}\geq e^{S_{t}}$. With this observation and using (48) we can now write
$e^{b}\leq{\mathbb{E}}_{\infty}[e^{S_{T_{\rm{L}}}}]\leq{\mathbb{E}}_{\infty}[R_{T_{\rm{L}}}]={\mathbb{E}}_{\infty}[T_{\rm{L}}]-1\leq{\mathbb{E}}_{\infty}[T_{\rm{L}}],$
which proves the desired inequality.
D-B Proof of Lemma IV.2
Similar to the proof of Lemma IV.1, we again define the Shiryaev-Roberts like statistic $\{R_{t}\}$. $\{R_{t}-t\}$ is again a martingale, thus ${\mathbb{E}}_{\infty}[T]-1={\mathbb{E}}_{\infty}[R_{T}]$ for the stopping time $T$.
Denote $L_{t}\overset{iid}{\sim}\rm{Lap}(\frac{4C}{\epsilon})$, $t\geq 1$. Then the detection statistics updated in (9), after exponentiation, can be equivalently written as
$e^{\tilde{S}_{t}}=e^{\mathcal{S}_{t}}e^{L_{t}}\leq R_{t}e^{L_{t}},t>1,$
where the inequality is due to $R_{t}\geq e^{\mathcal{S}_{t}}$ according to the proof of Lemma IV.1.
Write the randomized threshold as $\tilde{b}=b+L_{0}$ for $L_{0}\sim\rm{Lap}(\frac{2C}{\epsilon})$, then using (48) we can now write
$e^{b}\cdot\mathbb{E}[e^{L_{0}}]\leq{\mathbb{E}}_{\infty}[e^{\tilde{S}_{T}}]\leq{\mathbb{E}}_{\infty}[R_{T}]{\mathbb{E}}_{\infty}[e^{L_{T}}]\leq{\mathbb{E}}_{\infty}[T]{\mathbb{E}}_{\infty}[e^{L_{T}}].$
Note that for $L\sim\rm{Lap}(0,\beta)$, we have $\mathbb{E}[e^{L}]=\frac{1}{1-\beta^{2}}$ when $\beta<1$. Thus we have when $\epsilon>4C$,
${\mathbb{E}}_{\infty}[T]\geq\frac{\mathbb{E}[e^{L_{0}}]}{{\mathbb{E}}[e^{L_{T}}]}e^{b}=\frac{1-(4C/\epsilon)^{2}}{1-(2C/\epsilon)^{2}}e^{b},$
which proves the desired inequality.
D-C Proof of Theorem IV.5
Then we provide theoretical guarantees on the detection delay of the procedure (7) and (9). Recall $I_{0}=\frac{1}{2}(\log\frac{1-\zeta}{\zeta})p(1-2\zeta)({n\choose 2}-\sum_{i<j}\bm{\sigma}^{\text{pre}}_{i}\bm{\sigma}^{\text{pre}}_{j}\bm{\sigma}_{i}^{\text{post}}\bm{\sigma}_{j}^{\text{post}})={\mathbb{E}}_{1}[\ell({\mathbf{A}};\bm{\sigma}^{\text{pre}},\bm{\sigma}^{\text{post}})]$ is the KL divergence for original CMBs, and $\tilde{I}_{0}=\frac{1}{2}(\log\frac{1-\tilde{\zeta}}{\tilde{\zeta}})\tilde{p}(1-2\tilde{\zeta})({n\choose 2}-\sum_{i<j}\bm{\sigma}^{\text{pre}}_{i}\bm{\sigma}^{\text{pre}}_{j}\bm{\sigma}_{i}^{\text{post}}\bm{\sigma}_{j}^{\text{post}})={\mathbb{E}}_{1}[\ell(\tilde{\mathbf{A}};\bm{\sigma}^{\text{pre}},\bm{\sigma}^{\text{post}})]$ is the KL information number for the distribution after graph perturbation. Here ${\mathbb{E}}_{1}$ represents the expectation taken under the post-change regime (i.e., change-point $\nu=1$).
We first present the well-known information-theoretic lower bound on the detection delay for online change detection.
Theorem D.1 (Information-Theoretic Lower Bound on WADD [36]).
A lower bound on the WADD of any test that satisfies the average run length constraint $\mathbb{E}_{\infty}\left[\tau\right]\geq\gamma$ is
${\mathsf{WADD}}(\tau)\geq\frac{\log\gamma}{I}(1+o(1)),$
as $\gamma\to\infty$, where $I$ is the KL divergence of the post- and pre-change data distributions.
Based on Theorem D.1, when choosing $b=\log\gamma$, we have $\mathbb{E}_{\infty}[T_{\rm L}]\geq\gamma$, thus we have
${\mathsf{WADD}}(T_{\rm L})\geq\frac{\log\gamma}{\tilde{I}_{0}}(1+o(1))$ (49)
since by definition $\tilde{I}_{0}$ is the KL divergence between the post- and pre-change graph distributions after perturbation. We then list the following Lemma as a direct consequence of Theorem 1 of [39], observing that $\log\gamma$ dominates the numerator term under our assumption.
Lemma D.1 (WADD Upper Bound [39]).
@@
Under the assumptions of Theorem IV.5, we have the following upper bound for the worst-case performance of the detection procedure (8),
${\mathsf{WADD}}[T_{\rm L}]\leq\frac{\log\gamma}{\hat{I}_{0}}(1+o(1)),$ (50)
where $\hat{I}_{0}={\mathbb{E}}_{1}[\ell(\tilde{\mathbf{A}};\bm{\sigma}^{\text{pre}},\bm{\hat{\sigma}}_{t-1})]$.
In the following, we analyze the empirical quantities $\hat{I}_{0}$ and establish the asymptotic result. Following the proof of Theorem IV.1, we have under the post-change regime, the estimate $\bm{\hat{\sigma}}_{t}$ resulted from the most recent adjacency matrix satisfies, as $n\to\infty$,
$\operatorname{Pr}(\bm{\hat{\sigma}}_{t}=\bm{\sigma}^{\text{post}})\geq 1-n^{-\Omega(1)}.$
Then we have
$\hat{I}_{0}=\tilde{I}_{0}\Pr\{\bm{\hat{\sigma}}_{t}=\bm{\sigma}^{\text{post}}\}+{\mathbb{E}}[\hat{I}_{0}|\bm{\hat{\sigma}}_{t}\neq\bm{\sigma}^{\text{post}}]\Pr\{\bm{\hat{\sigma}}_{t}\neq\bm{\sigma}^{\text{post}}\}$
@@
note that $\mathbb{E}[\hat{I}_{0}|\bm{\hat{\sigma}}_{t}]=\mathbb{E}[\log\frac{\Pr(\tilde{\mathbf{A}}_{t};\hat{\bm{\sigma}}_{t})}{\Pr(\tilde{\mathbf{A}}_{t};\bm{\sigma}^{\text{pre}})}]=\mathbb{E}[\log\frac{\Pr(\tilde{\mathbf{A}}_{t};\hat{\bm{\sigma}}_{t})}{\Pr(\tilde{\mathbf{A}}_{t};\bm{\sigma}^{\text{post}})}+\log\frac{\Pr(\tilde{\mathbf{A}}_{t};\bm{\sigma}^{\text{post}})}{\Pr(\tilde{\mathbf{A}}_{t};\bm{\sigma}^{\text{pre}})}]\in[0,\tilde{I}_{0}]$. Therefore,
$\frac{\hat{I}_{0}}{\tilde{I}_{0}}$ $=\frac{\tilde{I}_{0}\Pr\{\bm{\hat{\sigma}}_{t}=\bm{\sigma}^{\text{post}}\}+{\mathbb{E}}[\hat{I}_{0}|\bm{\hat{\sigma}}_{t}\neq\bm{\sigma}^{\text{post}}]\Pr\{\bm{\hat{\sigma}}_{t}\neq\bm{\sigma}^{\text{post}}\}}{\tilde{I}_{0}}$
$=1-O(n^{-c}),$ (51)
@@
for some constant $c$, and when combined with (50), we obtain the desired asymptotic result in Theorem IV.5.
D-D Proof of Theorem IV.6
Based on Theorem D.1, we have ${\mathsf{WADD}}(T_{\rm C})\geq\frac{\log\gamma}{I_{0}}(1+o(1))$. It remains to prove $\frac{\log\gamma}{I_{0}}(1+o(1))$ is also an upper bound for ${\mathsf{WADD}}(T_{\rm C})$. To show this, we modify the proof of Theorem 1 in [39] to incorporate the additional Laplace noise added to the detection statistics.
Define the process $\{U_{t},t=1,2,\ldots\}$ as $U_{t}=U_{t-1}+\log\frac{\Pr({\mathbf{A}}_{t};\bm{\hat{\sigma}}_{t-1})}{\Pr({\mathbf{A}}_{t};\bm{\sigma}^{\text{pre}})}$ with $U_{1}=0$, and the corresponding stopping time $T^{\prime}=\inf\{t>1:U_{t}+L_{t}\geq\tilde{b}\}$ for the threshold $\tilde{b}=b+L_{0}$. Obviously we have $U_{t}\leq\mathcal{S}_{t}$, thus $U_{t}+L_{t}\leq\mathcal{S}_{t}+L_{t}=\tilde{S}_{t}$ and $\mathbb{E}[T_{\rm C}]\leq\mathbb{E}[T^{\prime}]$. In the following, we derive an upper bound for $\mathbb{E}[T^{\prime}]$.
For simplicity let us further denote with $\ell_{t}=\log\frac{\Pr({\mathbf{A}}_{t};\bm{\hat{\sigma}}_{t-1})}{\Pr({\mathbf{A}}_{t};\bm{\sigma}^{\text{pre}})}$ and $\hat{I}_{0}:={\mathbb{E}}_{1}[\ell_{t}]$, then we observe that for $t>1$ we can write
${\mathbb{E}}[U_{T^{\prime}}]$ $={\mathbb{E}}[\sum_{t=2}^{T^{\prime}}\ell_{t}]={\mathbb{E}}[\sum_{t=2}^{T^{\prime}+1}\ell_{t}]-{\mathbb{E}}_{0}[\ell_{T^{\prime}+1}]$
$={\mathbb{E}}[\sum_{t=2}^{\infty}\ell_{t}\mathbb{I}\{T^{\prime}\geq t-1\}]-{\mathbb{E}}_{0}[\ell_{T^{\prime}+1}]$
$={\mathbb{E}}[\sum_{t=2}^{\infty}{\mathbb{E}}\left[\ell_{t}|\mathcal{F}_{t-2}\right]\mathbb{I}\{T^{\prime}\geq t-1\}]-{\mathbb{E}}_{0}[\ell_{T^{\prime}+1}]$
$=\hat{I}_{0}{\mathbb{E}}[T^{\prime}]-{\mathbb{E}}_{0}[\ell_{T^{\prime}+1}]\geq\hat{I}_{0}{\mathbb{E}}[T^{\prime}]-I_{0}.$
Then we conclude that
${\mathbb{E}}[T^{\prime}]\leq\frac{{\mathbb{E}}[U_{T^{\prime}}]+I_{0}}{\hat{I}_{0}}=\frac{{\mathbb{E}}[U_{T^{\prime}}+L_{T^{\prime}}-\tilde{b}]+{\mathbb{E}}[\tilde{b}-L_{T^{\prime}}]+I_{0}}{\hat{I}_{0}}.$ (52)
The next step involves the analysis of the expectation of the overshoot $U_{T^{\prime}}+L_{T^{\prime}}-\tilde{b}$. We borrow ideas from [28] and modify the proof in [39]. For any threshold $x>0$ define $T^{\prime}_{x}$ to be the corresponding stopping time
$T^{\prime}_{x}=\inf\{t>1:U_{t}+L_{t}\geq x\},$
and denote the overshoot function as $R_{x}=U_{T^{\prime}_{x}}+L_{T^{\prime}_{x}}-x$. Define a sequence of stopping times $\{\tau_{j}\}$ with $\tau_{0}=1$ and
$\tau_{j}=\inf\left\{t>\tau_{j-1}:\sum_{s=\tau_{j-1}+1}^{t}\ell_{s}+L_{t}>L_{\tau_{j-1}}\right\},~j\geq 1,$
and the corresponding ladder variables $z_{j}=\sum_{s=\tau_{j-1}+1}^{\tau_{j}}\ell_{s}+L_{\tau_{j}}-L_{\tau_{j-1}}>0$. We can now see that $U_{\tau_{j}}+L_{\tau_{j}}=\sum_{i=1}^{j}z_{i}$, in fact the sequence $\{U_{t}+L_{t}\}_{t\geq 1}$ increases only at the stopping times $\{\tau_{j}\}$. For any given threshold $\nu$, in order to stop at $T_{\nu}^{\prime}$ the statistic $U_{t}+L_{t}$ needs an increase at $T_{\nu}^{\prime}$, which means that there exists a random index $j_{\nu}$ such that $\tau_{j_{\nu}}=T_{\nu}^{\prime}$. Due to this fact, we can write $R_{\nu}=\sum_{j=1}^{j_{\nu}}z_{j}-\nu$. Following the same steps as in [28], we observe that $R_{x}$ is a piecewise linear function and all pieces having slope 1-1, thus we have
$\int_{0}^{\nu}R_{x}\,dx$ $=\int_{0}^{U_{T^{\prime}_{\nu}}+L_{T^{\prime}_{\nu}}}R_{x}\,dx-\int_{\nu}^{U_{T^{\prime}_{\nu}}+L_{T^{\prime}_{\nu}}}R_{x}\,dx$ (53)
$=\frac{1}{2}\left\{\sum_{j=1}^{j_{\nu}}z_{j}^{2}-R_{\nu}^{2}\right\}.$
By definition $0<z_{j}$ and $z_{j}=\sum_{t=\tau_{j-1}+1}^{\tau_{j}-1}\ell_{t}+\ell_{\tau_{j}}+L_{\tau_{j}}-L_{\tau_{j-1}}$ with $\sum_{t=\tau_{j-1}+1}^{\tau_{j}-1}\ell_{t}+L_{\tau_{j}-1}-L_{\tau_{j-1}}\leq 0$, then we have $0<z_{j}\leq\ell_{\tau_{j}}+L_{\tau_{j}}-L_{\tau_{j}-1}$, which implies $z_{j}^{2}\leq 2\ell_{\tau_{j}}^{2}+2(L_{\tau_{j}}-L_{\tau_{j}-1})^{2}\leq 2\sum_{t=\tau_{j-1}+1}^{\tau_{j}}\left[\ell_{t}^{2}+(L_{t}-L_{t-1})^{2}\right]$. Substituting in (53) yields
$\int_{0}^{\nu}R_{x}\,dx\leq\sum_{t=2}^{T^{\prime}+1}[\ell_{t}^{2}+(L_{t}-L_{t-1})^{2}]-\frac{1}{2}R_{\nu}^{2}.$
If we take the expectation of above expression and use Jensen’s inequality on the last term we obtain
$0\leq\int_{0}^{\nu}{\mathbb{E}}[R_{x}]\,dx\leq{\mathbb{E}}[\sum_{t=2}^{T^{\prime}+1}[\ell_{t}^{2}+(L_{t}-L_{t-1})^{2}]]-\frac{1}{2}({\mathbb{E}}[R_{\nu}])^{2}$
$=\frac{\hat{J}_{0}}{\hat{I}_{0}}{\mathbb{E}}[\sum_{t=2}^{T^{\prime}+1}\ell_{t}]-\frac{1}{2}({\mathbb{E}}[R_{\nu}])^{2}$
$=\frac{\hat{J}_{0}}{\hat{I}_{0}}{\mathbb{E}}\left[U_{T^{\prime}}+\ell_{T^{\prime}+1}\right]-\frac{1}{2}({\mathbb{E}}[R_{\nu}])^{2}$
$=\frac{\hat{J}_{0}}{\hat{I}_{0}}{\mathbb{E}}\left[R_{\nu}+\nu-L_{T^{\prime}}+\ell_{T^{\prime}+1}\right]-\frac{1}{2}({\mathbb{E}}[R_{\nu}])^{2}$
$\leq\frac{\hat{J}_{0}}{\hat{I}_{0}}\left\{{\mathbb{E}}[R_{\nu}]+\nu+I_{0}\right\}-\frac{1}{2}({\mathbb{E}}[R_{\nu}])^{2},$
where $\hat{J}_{0}=\mathbb{E}[\ell_{2}^{2}+(L_{2}-L_{1})^{2}]=\mathbb{E}[\ell_{2}^{2}]+64(\frac{C}{\epsilon})^{2}$; the first equality is true because ${\mathbb{E}}[\sum_{t=2}^{T^{\prime}+1}[\ell_{t}^{2}+(L_{t}-L_{t-1})^{2}]]=\hat{J}_{0}{\mathbb{E}}[T^{\prime}]$ and ${\mathbb{E}}[\sum_{t=2}^{T^{\prime}+1}\ell_{t}]=\hat{I}_{0}{\mathbb{E}}[T^{\prime}]$. From the nonnegativity of the integral we have $\frac{1}{2}({\mathbb{E}}[R_{\nu}])^{2}\leq\frac{\hat{J}_{0}}{\hat{I}_{0}}\left\{{\mathbb{E}}[R_{\nu}]+\nu+I_{0}\right\}$, from which we conclude that ${\mathbb{E}}[R_{\nu}]\leq 2\frac{\hat{J}_{0}}{\hat{I}_{0}}+\big(2\frac{\hat{J}_{0}}{\hat{I}_{0}}\nu\big)^{1/2}+\big(2\frac{\hat{J}_{0}}{\hat{I}_{0}}I_{0}\big)^{1/2}$. Substitute $\nu=\tilde{b}=b+L_{0}$ as a random threshold, we obtain
$\mathbb{E}[R_{\tilde{b}}$] $\leq 2\frac{\hat{J}_{0}}{\hat{I}_{0}}+{\mathbb{E}}[\big(2\frac{\hat{J}_{0}}{\hat{I}_{0}}\nu\big)^{1/2}]+\big(2\frac{\hat{J}_{0}}{\hat{I}_{0}}I_{0}\big)^{1/2}$
$\leq 2\frac{\hat{J}_{0}}{\hat{I}_{0}}+\big(2\frac{\hat{J}_{0}}{\hat{I}_{0}}\big)^{1/2}(\sqrt{b}+\frac{\sqrt{4C\pi/\epsilon}}{2})+\big(2\frac{\hat{J}_{0}}{\hat{I}_{0}}I_{0}\big)^{1/2}.$
Substitute above into (52), and plug in $b=\log\gamma+\log\frac{1-(2C/\epsilon)^{2}}{1-(4C/\epsilon)^{2}}$, we obtain
$\mathbb{E}[T^{\prime}]$ $\leq\frac{\log\gamma+2\frac{\hat{J}_{0}}{\hat{I}_{0}}+\big(2\frac{\hat{J}_{0}}{\hat{I}_{0}}\big)^{1/2}(\sqrt{b}+\frac{\sqrt{4C\pi/\epsilon}}{2})}{\hat{I}_{0}}$
$\hskip 7.22743pt+\frac{\big(2\frac{\hat{J}_{0}}{\hat{I}_{0}}I_{0}\big)^{1/2}+\log\frac{1-(2C/\epsilon)^{2}}{1-(4C/\epsilon)^{2}}+I_{0}}{\hat{I}_{0}}.$
We note that $\log\frac{1-(2C/\epsilon)^{2}}{1-(4C/\epsilon)^{2}}=o(\log\gamma)$ and $(\frac{\hat{J}_{0}}{\hat{I}_{0}})^{1/2}\sqrt{4C\pi/\epsilon}=o(\log\gamma)$. Similar to the proof of Theorem IV.5, we also have $\frac{\hat{I}_{0}}{I_{0}}=1-O(n^{-c^{\prime}})$ for some constant $c^{\prime}$ under exact recovery condition. Then the above upper bound can be written as $\mathbb{E}[T^{\prime}]\leq\frac{\log\gamma}{I_{0}}(1+o(1))$. This completes the proof.
Appendix E Proofs for Section IV-C
E-A Proof of Lemma IV.3